# Combining straight-line and map-based distances to investigate the connection between proximity to healthy foods and disease

### v1

@@ 3 Simulation Studies
Using data generated to mimic the Piedmont Triad data analyzed in Section 4, we demonstrate how (i) using error-prone, straight-line food access measures can lead to bias, but (ii) incorporating more accurate, map-based food access measures for even a small subset of neighborhoods can correct that bias. Further, imputing map-based food access for neighborhoods where it was missing can offer better efficiency than ignoring those neighborhoods. R scripts to reproduce all simulations, tables, and figures are available on GitHub at https://github.com/sarahlotspeich/food_access_imputation.
@@ 3.1 Setup and Data Generation
The model in Equation (1) was of interest, which captured the association between the neighborhood-level prevalence of the outcome ($Y/Pop$) and map-based proximity to healthy food stores ($X$) using Poisson regression. For comparison, this model was fit to the unqueried data (i.e., using $X^{*}$ for all neighborhoods), the fully queried data (i.e., using $X$ for all neighborhoods), and the partially queried data (i.e., using $X$ and $X^{*}$ for some neighborhoods and just $X^{*}$ for the rest). The gold standard analysis used the fully queried data to fit the model using $Y$, $Pop$, and $X$ from all $N$ neighborhoods. The naive analysis used the unqueried data to fit the model using $Y$, $Pop$, and $X^{*}$ (instead of $X$) from all $N$ neighborhoods. The complete-case analysis used the partially queried data to fit the model using $Y$, $Pop$, and $X$ from only the subset of $n$ queried neighborhoods. The imputation analysis used the partially queried and unqueried data together to fit the model using $Y$, $Pop$, and multiply imputed $\widetilde{X}$ (instead of $X$) from all $N$ neighborhoods $B=20$ times following the process detailed in Section 2.3.
Empirical relative bias and standard errors for the estimated log prevalence ratio $\hat{\beta}_{1}$ for map-based proximity $X$ are reported for the four analysis approaches to assess their validity for estimation. For imputation, the average standard error estimates are compared to the empirical standard errors to assess its validity for inference. Among unbiased approaches, relative efficiency to the gold standard analysis is reported to compare statistical precision. Relative efficiency is calculated as the ratio of the empirical variances of the gold standard to the other approaches, with ratios closer to one indicating that more efficiency was recovered.
@@ 3.2 Additive Errors in Straight-Line Proximity
Simulation results under increasingly severe additive errors in straight-line proximity to healthy foods $X^{*}$ are summarized in Table 1. The error magnitude (i.e., severity) was varied by considering different standard deviations $\sigma_{U}\in\{0.1,0.2,0.4,0.8,1\}$ for $U$. For all choices of $\sigma_{U}$, straight-line proximity $X^{*}$ remained less than or equal to map-basedp proximity $X$. With $\hat{\sigma}_{U}=0.8$ in the Piedmont Triad data (Section 4.4), more and less severe settings were considered. As expected, the relative bias of the naive analysis grew as $\sigma_{U}$ increased and persisted regardless of $N$. In most cases, the complete-case analysis offered reduced bias over the naive, particularly under more severe errors. Imputation saw lower bias than the naive and complete-case analyses in almost all settings. Even with the smallest sample size and most severe errors, imputation was $<1\%$ biased.
Table 1: Simulation results under varied additive errors in straight-line proximity to healthy foods, as controlled by the standard deviation $\sigma_{U}$ of the errors $U$.
@@
The standard error estimator for imputation performed well, closely approximating the empirical standard error on average. Empirical coverage probabilities for its 95% confidence intervals were close to $0.95$. Given the small scale of the true $\beta_{1}=0.01$, the standard errors for all approaches were small. Still, the imputation estimator recovered $50-99\%$ of the efficiency of the gold standard analysis, while the complete case analysis only recovered up to $11\%$. With either $N$, the imputation estimator recovered more efficiency with smaller $\sigma_{U}$, when the relationship between $X$ and $X^{*}$ was more informative. Similar observations hold when the error mean $\mu_{U}$ was varied instead (Supplemental Table S1). Thus, imputation performs well under errors in $X^{*}$ that underestimated $X$ by a bigger margin (driven by larger $\mu_{U}$) in addition to errors $X^{*}$ that were noisier about $X$ (due to larger $\sigma_{U}$).
Simulation results with a different proportion $q\in\{0.1,0.25,0.5,0.75\}$ of queried neighborhoods with non-missing $X$ can be found in Table 2. All metrics for the naive analysis, which ignores the queried data, were effectively unchanged across these settings. For any sample size $N$, the standard errors and, to a lesser extent, the relative bias of the complete-case analysis and imputation estimators decreased as $q$ increased (i.e., when there was less missing data). While the efficiency gains of imputation over the complete-case analysis were largest when $q$ was small, they remained sizeable (relative efficiency $18-24\%$ larger) even when $q$ was large.
Table 2: Simulation results with more neighborhoods queried to obtain map-based proximity to healthy foods, as controlled by the proportion $q=n/N$.
@@
Table 3 summarizes simulations where the disease prevalence and prevalence ratio for map-based food access $X$ were varied. Average disease prevalences (hereafter, simply “prevalences”) of $7\%$, $11\%$, and $35\%$ were considered by varying $\beta_{0}\in\{\log(0.07),\log(0.11),\log(0.35)\}$, respectively, and prevalence ratios $\exp(\beta_{1})\in\{0.95,0.99,1,1.01,1.05\}$ were included. At the null (i.e., $\beta_{1}=0$), all methods were unbiased, regardless of the prevalence. Beyond that setting, the bias of the naive analysis increased in magnitude as $\beta_{1}$ got farther from zero but was comparable across prevalences. The complete-case and imputation estimators were both virtually unbiased, and the confidence intervals for the imputation estimator achieved proper coverage in all settings. The complete-case analysis only recovered $8-9\%$ of the efficiency of the gold standard. Meanwhile, imputation recovered up to $70\%$ at the null but less as $\beta_{1}$ moved away from zero.
@@ 3.3 Multiplicative Errors in Straight-Line Proximity
Error-prone $X^{*}$ was also simulated from a multiplicative error model such that $X^{*}=WX$. Random errors $W$ were generated from a truncated Normal distribution with mean $=0.7$, standard deviation $=\sigma_{W}$, and upper bound $=1$. (Setting the maximum to $1$ simulated straight-line proximity to be strictly less than or equal to map-based, as expected.) Different standard deviations $\sigma_{W}\in\{0.1,0.15,0.2\}$ were considered for $W$, where smaller values led to less severe errors in $X^{*}$. With $\hat{\sigma}_{W}=0.15$ in the Piedmont Triad data, more and less severe settings were considered. All other variables were simulated following Section 3.1.
Simulation results under increasingly severe multiplicative errors in straight-line proximity to healthy foods are described in Table 4. The naive analysis was considerably more biased in the multiplicative error settings considered than under the additive ones in Section 3.2 (as high as $37\%$ versus $8\%$). The complete-case analysis offered low bias ($\leq 1\%$) but low efficiency ($10\%$ or less recovered) in all settings. The imputation estimator offered similarly low bias ($<2\%$) with much higher efficiency ($39-62\%$ recovered). Its standard error estimator and confidence intervals remained valid in all settings. As with additive errors, the imputation estimator recovered more efficiency in the least severe settings (i.e., with smaller $\sigma_{W}$) in small or large samples.
Table 4: Simulation results under varied multiplicative errors in straight-line proximity to healthy foods, as controlled by the standard deviation $\sigma_{W}$ of the errors $W$.
@@ 4 Application to the Piedmont Triad, North Carolina
The Piedmont Triad, located in the northwestern region of North Carolina, comprises twelve counties. Beyond two medium-sized cities (Greensboro and Winston–Salem), much of the area is rural, with a median population density for neighborhoods (census tracts) across the region of $843$ people per square mile (minimum $=33$, maximum $=6681$). The region includes one neighborhood with zero population, which contains only the Piedmont Triad International Airport and is excluded from all analyses. The $N=387$ neighborhoods in the Piedmont Triad are described in Supplemental Table S2. The number of neighborhoods per county varied from as few as six in the rural counties to as many as $118$ in the urban ones. According to the 2010 U.S. Census, the median population per neighborhood was $4095$ people (IQR $=[3901,5282]$). Land area per census tract was $5$ square miles, on average (IQR $=[2,19]$).
The socioeconomic factors vary substantially between neighborhoods in this region. A brief socioeconomic profile, based on the 2015 American Community Survey (ACS), is illustrated in Supplemental Figure S4. The median household income was \$53,750, on average, but was as low as \$13,447 and as high as \$176,875. The percentage of neighborhood households with income below the poverty line varies widely and ranges from $0\%$ and $73\%$. Most neighborhoods had low proportions of households receiving the Supplemental Nutrition Assistance Program (SNAP) (median $=13\%$). Still, there were some neighborhoods with as many as $75\%$ who were receiving this type of government assistance. Most workers ($\geq 45\%$) in nearly all neighborhoods drive alone to work; still, a neighborhood where $97\%$ of workers drive alone (the maximum) would be very different than one where $45\%$ do (the minimum). The proportion of people with health insurance (public or private) varied some (between $65\%$ and $100\%$), although not as much as other variables like the proportion of people $25$ years or older who completed at least some college (between $26\%$ and $96\%$) and the others mentioned thus far.
@@ 4.1 Motivation for Analysis
Preliminary data using straight-line proximity to healthy foods suggest that many neighborhoods in the Piedmont Triad, NC, have few options close to home (Figure 1). The actual number of neighborhoods with poor access to healthy foods is likely even higher, since map-based proximity has to be as bad as or worse than straight-line proximity. This concern is echoed by data reported from the Economic Research Service , U.S. Department of Agriculture (USDA)(2019) (ERS), wherein $46\%$ and $67\%$ of neighborhoods in the Piedmont Triad were found to have high proportions of people living more than $1$ mile and $0.5$ mile, respectively, from their nearest supermarket. Many neighborhoods without access are in Forsyth and Guilford County, putting these counties among the least accessible in the state. Interestingly, Forsyth and Guilford are both relatively urban counties, whereas one might expect lower access in more expansive, rural areas. However, beyond these cities’ largest cities (Winston-Salem and Greensboro, respectively), there are many rural neighborhoods that could help explain this result.
Figure 1: Choropleth map of food access, as measured by proximity to healthy foods for each neighborhood (census tract) in the Piedmont Triad, North Carolina, according to the data used for the naive, gold standard, imputation, and complete-case analyses.
National studies in the U.S. have previously found links between diabetes and obesity and the availability of healthy foods. Ahern et al. (2011) found that lower county-level prevalence of type 2 diabetes was associated with having more “healthy” food options (full-service restaurants, grocery stores, direct farm sales). Kanchi et al. (2021) found that higher relative densities of supermarkets to other retail food outlets were associated with lessened risk of incident type 2 diabetes in a national cohort study of U.S. veterans. Ahern et al. (2011) also found full-service restaurants and direct farm sales to be associated with lower county-level prevalence of obesity, while grocery stores were associated with higher rates. However, other studies have found conflicting evidence on these connections between food access and health (e.g., White, 2007), so it is critical to re-investigate these relationships for the Piedmont Triad alone.
@@ 4.2 Data Collection
Publicly available data were combined from three sources for the analysis using open-source tools in R (R Core Team, 2023). The tidycensus package was used to extract ACS data and create maps (Walker and Herman, 2024). Links to the datasets and all R code are available on GitHub at https://github.com/sarahlotspeich/food_access_imputation.
Diabetes and obesity prevalence estimates were taken from the U.S. Centers for Disease Control and Prevention’s 2022 PLACES dataset (Centers for Disease Control and Prevention, 2022). Like other administrative datasets, PLACES contains small area estimates (SAEs) of prevalence, which were obtained via statistical modeling using individual- and area-level information (e.g., age, education, poverty) as predictors (Centers for Disease Control and Prevention, 2023). The PLACES data were available at the census tract level, which was adopted as the “neighborhood” unit for analysis.
Store locations and types were obtained from the U.S. Department of Agriculture’s 2022 Historical SNAP Retail Locator Data (United States Department of Agriculture, 2022) to calculate neighborhoods’ proximity to healthy foods. The analysis used $M=701$ NC farmers’ markets, grocery stores (any size), specialties (baker, fruits/vegetables, meat/poultry, seafood), and supermarkets/stores as “healthy food stores.” A map of stores by type can be found in Supplemental Figure S5.
With a moderate number of neighborhoods in the Piedmont Triad ($N=387$ census tracts), we were actually able to establish the gold standard here, which would not always be possible in practice. Using Google Maps and the Haversine formula for distance calculations, we collected map-based and straight-line proximity, respectively, to healthy foods for the entire region. Then, map-based proximity $X$ from these fully queried data was used to fit the gold standard analysis. The naive analysis used straight-line proximity $X^{*}$ from the unqueried data instead.
For the complete-case and imputation analyses, an artificial partially queried dataset was also constructed, wherein only $n=48$ census tracts ($q=0.12$) had map-based $X$ available. To ensure geographic diversity of this queried subset, four neighborhoods were chosen from each county to be treated as queried and have map-based proximity $X$ available (Supplemental Figure S6). Many other sampling designs are possible, and targeted strategies are a promising direction for future work.

### v2

@@ 3 Simulation Studies
Using data generated to mimic the Piedmont Triad data analyzed in Section 4, we demonstrate how (i) using error-prone, straight-line food access measures can lead to bias but (ii) incorporating more accurate, map-based food access measures for even a small subset of neighborhoods can correct that bias. Further, imputing map-based food access for neighborhoods where it was missing can offer better efficiency than ignoring those neighborhoods. R scripts to reproduce all simulations, tables, and figures are available on GitHub at https://github.com/sarahlotspeich/food_access_imputation.
@@ 3.1 Setup and data generation
The model in Equation (1) was of interest, which captured the association between the neighborhood-level prevalence of the outcome ($Y/Pop$) and map-based proximity to healthy food stores ($X$) using Poisson regression. For comparison, this model was fit to the unqueried data (i.e., using $X^{*}$ for all neighborhoods), the fully queried data (i.e., using $X$ for all neighborhoods), and the partially queried data (i.e., using $X$ for a subset of neighborhoods). The gold standard analysis used the fully queried data to fit the model using $Y$, $Pop$, and $X$ from all $N$ neighborhoods. The naive analysis used the unqueried data to fit the model using $Y$, $Pop$, and $X^{*}$ (instead of $X$) from all $N$ neighborhoods. The complete case analysis used the partially queried data to fit the model using $Y$, $Pop$, and $X$ from only the subset of $n$ neighborhoods. The imputation analysis used the partially queried and unqueried data together to fit the model using $Y$, $Pop$, and multiply imputed $\widetilde{X}$ (instead of $X$) from all $N$ neighborhoods $B=20$ times following the process detailed in Section 2.3.
Empirical relative bias and standard errors for the estimated log prevalence ratio $\hat{\beta}_{1}$ for map-based proximity $X$ are reported for the four analysis approaches to assess their validity for estimation. For imputation, the average standard error estimates are compared to the empirical standard errors to assess their validity for inference. Among unbiased approaches, relative efficiency to the gold standard analysis is reported to compare statistical precision. Relative efficiency is calculated as the ratio of the empirical variances of the gold standard to the other approaches, with ratios closer to one indicating that more efficiency was recovered.
@@ 3.2 Additive errors in straight-line proximity
Simulation results under increasingly severe additive errors in straight-line proximity to healthy foods $X^{*}$ are summarized in Table 1. The error magnitude (i.e., severity) was varied by considering different standard deviations $\sigma_{U}\in\{0.1,0.2,0.4,0.8,1\}$ for $U$. For all choices of $\sigma_{U}$, straight-line proximity $X^{*}$ remained less than or equal to map-based proximity $X$. With $\hat{\sigma}_{U}=0.8$ in the Piedmont Triad data (Section 4.4), our simulations considered more and less severe errors than were in the real data. As expected, the relative bias of the naive analysis grew as $\sigma_{U}$ increased (from $3\%$ to $6\%$) and persisted regardless of $N$. In all cases, the complete case and imputation analyses offered reduced bias over the naive, particularly under more severe errors. Even with the smallest sample size and most severe errors, imputation was $<2\%$ biased.
Table 1: Simulation results under increasingly severe additive errors in straight-line proximity to healthy foods, as controlled by the standard deviation $\sigma_{U}$ of the errors $U$. The mean $\mu_{U}=-0.7$ of the errors was fixed.
@@
The standard error estimator for imputation performed well, closely approximating the empirical standard error on average. In most settings, empirical coverage probabilities for its 95% confidence intervals were close to $0.95$. With the worst errors ($\sigma_{U}=0.8$ or $1$), the confidence intervals were conservative (i.e., with higher-than-nominal coverage). Since $X$ was simulated from a gamma distribution and the error $U$ added to it from a truncated normal, the normal linear regression imputation model is technically misspecified here, which could explain the inflated coverage. [51] Either the Robins and Wang estimator[52] or a bootstrap approach [53] might be adopted to obtain narrower confidence intervals.
Given the small scale of the true $\beta_{1}=0.01$, the standard errors for all approaches were small. Still, the imputation estimator recovered $57-93\%$ of the efficiency of the gold standard analysis, while the complete case analysis only recovered up to $10\%$. With either $N$, the imputation estimator recovered more efficiency with smaller $\sigma_{U}$, when the relationship between $X$ and $X^{*}$ was more informative. Similar observations hold when the error mean $\mu_{U}$ was varied instead (Supplemental Table S1). Thus, imputation performs well under errors in $X^{*}$ that underestimated $X$ by a bigger margin (driven by larger $\mu_{U}$) in addition to errors $X^{*}$ that were noisier about $X$ (due to larger $\sigma_{U}$).
Simulation results with a different proportion $q\in\{0.1,0.25,0.5,0.75\}$ of queried neighborhoods with non-missing $X$ can be found in Table 2. All metrics for the naive analysis, which ignores the queried data, were effectively unchanged across these settings. For any sample size $N$, the standard errors and, to a lesser extent, the relative bias of the complete case analysis and imputation estimators decreased as $q$ increased (i.e., when there was less missing data). While the efficiency gains of imputation over the complete case analysis were largest when $q$ was small (relative efficiency $56\%$ larger), they remained sizable (relative efficiency $21-22\%$ larger) even when $q$ was large.
Table 2: Simulation results with more neighborhoods queried to obtain map-based proximity to healthy foods, as controlled by the proportion $q=n/N$. The additive errors $U$ in straight-line proximity were generated with mean $\mu_{U}=-0.7$ and standard deviation $\sigma_{U}=0.8$.
@@
Table 3 summarizes simulations where the disease prevalence and prevalence ratio for map-based food access $X$ were varied. Average disease prevalences (hereafter, simply “prevalences”) of $7\%$, $11\%$, and $33\%$ were considered by varying $\beta_{0}\in\{-2.7,-2.2,-1.1\}$, respectively, and prevalence ratios of $\exp(\beta_{1})\in\{0.95,0.99,1,1.01,1.05\}$ were included. At the null (i.e., $\beta_{1}=0$), all methods were unbiased, regardless of the prevalence. Beyond that setting, the relative bias of the naive analysis increased as $\beta_{1}$ got farther from zero but was comparable across prevalences. The complete case and imputation estimators were both virtually unbiased ($<2\%$), and the confidence intervals for the imputation estimator achieved proper coverage in all settings. The complete case analysis only recovered $7-10\%$ of the efficiency of the gold standard. Meanwhile, imputation recovered up to $72\%$ at the null but less as $\beta_{1}$ moved away from zero.
@@ 3.3 Multiplicative errors in straight-line proximity
Error-prone $X^{*}$ was also simulated from a multiplicative error model such that $X^{*}=WX$. Random errors $W$ were generated from a truncated Normal distribution with mean $=\mu_{W}$, standard deviation $=\sigma_{W}$, lower bound $=0$, and upper bound $=1$. (Setting the maximum to $1$ simulated straight-line proximity to be strictly less than or equal to map-based, as expected.) Fixing $\mu_{W}=0.7$, different standard deviations $\sigma_{W}\in\{0.1,0.15,0.2\}$ were considered for $W$, where smaller values were expected to yield less severe errors in $X^{*}$. However, since the error mean of $0.7$ was close to the upper bound of $1$, larger values of $\sigma_{W}$ could actually lead to some more severe errors (with $W$ closer to 00) and some less severe errors (with $W$ closer to $1$). With $\hat{\sigma}_{W}=0.15$ in the Piedmont Triad data, settings with more and less error variability were again considered. All other variables were simulated following Section 3.1.
Simulation results under varied standard deviations for the multiplicative error in straight-line proximity to healthy foods are described in Table 4. The naive analysis was considerably more biased in the multiplicative error settings considered than under the additive ones in Section 3.2 (as high as $38\%$ versus $6\%$). The complete case analysis offered low bias ($<3\%$) but low efficiency ($10\%$ or less recovered) in all settings. The imputation estimator offered similarly low bias ($<2\%$) with much higher efficiency ($34-62\%$ recovered). Its standard error estimator and confidence intervals remained valid in all settings.
Table 4: Simulation results under varied multiplicative errors in straight-line proximity to healthy foods, as controlled by the standard deviation $\sigma_{W}$ of the errors $W$. The mean $\mu_{W}=0.7$ of the errors was fixed.
@@
As with additive errors, the imputation estimator recovered more efficiency in the settings with smaller $\sigma_{W}$. However, whereas larger $\sigma_{U}$ led to larger bias for the naive analysis with additive errors, larger $\sigma_{W}$ led to smaller bias for the naive analysis with multiplicative errors. Intuitively, this result is likely because the mean $\mu_{W}=0.7$ was already close to the upper bound of $1$. Thus, while increasing $\sigma_{W}$ pushed some simulated errors $W$ toward 00 (more severe), it also led to more errors near $1$ (less severe). In additional simulations, we varied the multiplicative error mean $\mu_{W}\in\{0.3,0.5,0.7\}$ but held the standard deviation $\sigma_{W}=0.15$ fixed. We saw that pushing the mean closer to the lower bound of 00 instead led to larger bias for the naive analysis, as we incurred more small values of $W$ and thus more severe errors (Supplemental Table S2).
@@ 4 Application to the Piedmont Triad, North Carolina
The Piedmont Triad, located in the northwestern region of North Carolina, comprises twelve counties (Supplemental Figure S4). Beyond two medium-sized cities (Greensboro and Winston–Salem), much of the area is rural, with a median population density for neighborhoods (census tracts) across the region of $843$ people per square mile (minimum $=33$, maximum $=6681$). The region includes one neighborhood with zero population, which contains only the Piedmont Triad International Airport and was excluded from all analyses. The remaining $N=387$ neighborhoods in the Piedmont Triad are described in Supplemental Table S3. The number of neighborhoods per county varied from as few as six in the rural counties to as many as $118$ in the urban ones. According to the 2010 U.S. Census, the median population per neighborhood was $4095$ people (IQR $=[3901,5282]$). Land area per neighborhood was $5$ square miles, on average (IQR $=[2,19]$).
The socioeconomic factors vary substantially between neighborhoods in this region. A brief socioeconomic profile based on the 2015 American Community Survey (ACS) is illustrated in Supplemental Figure S5. The median household income was \$53,750, on average, but was as low as \$13,447 and as high as \$176,875. The percentage of neighborhood households with income below the poverty line varies widely and ranges from $0\%$ to $73\%$. Most neighborhoods had low proportions of households receiving the Supplemental Nutrition Assistance Program (SNAP) (median $=13\%$). Still, there were some neighborhoods with as many as $75\%$ who were receiving this type of government assistance. Most workers ($\geq 45\%$) in nearly all neighborhoods drive alone to work; still, a neighborhood where $97\%$ of workers drive alone (the maximum) would be very different than one where $45\%$ do (the minimum). The proportion of people with health insurance (public or private) varied some (between $65\%$ and $100\%$), although not as much as other variables like the proportion of people $25$ years or older who completed at least some college (between $26\%$ and $96\%$) and the others mentioned thus far.
The average White and Black populations (as percents) per neighborhood in the Piedmont Triad were similar to the state averages of (78% versus 75% and 14% versus 15%). Like the rest of the state, the Piedmont Triad had extremely low percentages of Asian, American Indian or Alaska Native, or Native Hawaiian and Other Pacific Islander populations (all $\approx 0\%$, on average). Still, there was wide variability between neighborhoods (Supplemental Figure S6). In particular, some neighborhoods were more than 75% Black, and these neighborhoods were located primarily in either the urban areas surrounding the cities of Winston-Salem and Greensboro or in the rural county in the northeastern corner of the Piedmont Triad. While the average percent Asian population across the region was small, there were some neighborhoods with more than 20% Asian population, as well.
@@ 4.1 Motivation for analysis
Preliminary data using straight-line proximity to healthy foods suggested that many neighborhoods in the Piedmont Triad, NC, have few options close to home (Figure 1). We knew that the actual number of neighborhoods with poor access to healthy foods was likely even higher since map-based proximity had to be as bad as or worse than straight-line proximity. This concern was echoed by data reported in the U.S. Department of Agriculture’s Food Research Atlas, [54] wherein $46\%$ and $67\%$ of neighborhoods in the Piedmont Triad were found to have high proportions of people living more than $1$ mile and $0.5$ mile, respectively, from their nearest supermarket. Many neighborhoods without access were in Forsyth and Guilford County, putting these counties among the least accessible in the state. Interestingly, Forsyth and Guilford are both relatively urban counties, whereas one might expect lower access in more expansive, rural areas. However, beyond these counties’ largest cities (Winston-Salem and Greensboro, respectively), there are many rural neighborhoods that could help explain this result.
Figure 1: Choropleth map of food access, as measured by proximity to healthy foods for each neighborhood (census tract) in the Piedmont Triad, North Carolina, according to the data used for the naive, gold standard, imputation, and complete case analyses. One census tract was excluded because it had a population of zero.
National studies in the U.S. have previously found links between diabetes and obesity and the availability of healthy foods. Ahern et al. found that lower county-level prevalence of type 2 diabetes was associated with having more “healthy” food options (full-service restaurants, grocery stores, direct farm sales). [8] Kanchi et al. found that higher relative densities of supermarkets to other retail food outlets were associated with lessened risk of incident type 2 diabetes in a national cohort study of U.S. veterans. [10] Ahern et al. also found full-service restaurants and direct farm sales to be associated with lower county-level prevalence of obesity, while grocery stores were associated with higher rates. [8] However, other studies have found conflicting evidence on these connections between food access and health, [55] so it is critical to re-investigate these relationships for the Piedmont Triad alone.
@@ 4.2 Data collection
Publicly available data were combined from three sources for the analysis using open-source tools in R. [21] The tidycensus package was used to extract ACS data and create maps. [56] Links to the datasets and all R code are available on GitHub at https://github.com/sarahlotspeich/food_access_imputation.
Diabetes and obesity prevalence estimates were taken from the U.S. Centers for Disease Control and Prevention’s 2022 PLACES dataset. [57] Like other administrative datasets, PLACES contains small area estimates (SAEs) of prevalence, which were obtained via statistical modeling using individual- and area-level information (e.g., age, race, education, poverty) as predictors. [58] The PLACES data were available at the census tract level, which was adopted as the “neighborhood” unit for analysis.
Store locations and types were obtained from the U.S. Department of Agriculture’s 2022 Historical SNAP Retail Locator Data to calculate neighborhoods’ proximity to healthy foods. [59] The analysis used $M=701$ NC farmers’ markets, grocery stores (any size), specialties (baker, fruits/vegetables, meat/poultry, seafood), and supermarkets/stores as “healthy food stores.” A map of stores by type can be found in Supplemental Figure S7. To adjust for the urbanicity of the census tracts in analyses, the 2010 Rural-Urban Commuting Area (RUCA) codes were also extracted from the U.S. Department of Agriculture. [60]
With a moderate number of neighborhoods in the Piedmont Triad ($N=387$ census tracts), we were able to establish the gold standard here, which is not always possible in practice. Using Google Maps and the Haversine formula for distance calculations, we collected map-based and straight-line proximity, respectively, to healthy foods for the entire region. Then, map-based proximity $X$ from these fully queried data was used to fit the gold standard analysis. The naive analysis used straight-line proximity $X^{*}$ from the unqueried data instead.
For the complete case and imputation analyses, an artificial partially queried dataset was also constructed, wherein only $n=48$ census tracts ($q=0.12$) had map-based $X$ available. To ensure the geographic diversity of this queried subset, four neighborhoods were chosen from each county at random to be treated as queried and have map-based proximity $X$ available (Supplemental Figure S8). This design targets equal representation from counties made up of (i) a large number of small, urban census tracts and (ii) a small number of large, rural census tracts. The former can be found in the center of the Piedmont Triad, where large cities like Winston-Salem and Greensboro are located, while the latter are found along the border of the region. Many other sampling designs are possible, and analysis-specific, targeted strategies are a promising direction for future work.